# Leveraging Machine Learning for High-Dimensional Option Pricing within the Uncertain Volatility Model

### v1

@@ 5.3 Path dependent option: the Call Sharpe option
Table 9: Call Sharpe. The column labeled “GL” presents the results from [6] for the same option, while the column labeled “BM” shows the benchmark computed by using a PDE approach. Values in brackets indicate the computational times, measured in seconds.
@@ 6 Conclusions
In this study, we demonstrated the efficacy of Machine Learning in enhancing the Uncertain Volatility Model for high-dimensional option pricing. Our findings suggest that Gaussian Process Regression algorithm can effectively handle the complexity and variability inherent in financial markets, providing more accurate and reliable pricing under the UVM framework. The integration of Machine Learning not only improves the precision of the optimal volatility and correlation estimates but also optimizes computational efficiency, making it a valuable tool for traders and risk managers. Furthermore, our approach offers a robust method for superhedging strategies, ensuring non-negative profit and loss outcomes even in highly volatile market conditions. Future research can extend this work by exploring other ML techniques and their applications in different financial models, further bridging the gap between theoretical advancements and practical implementations in the field of quantitative finance.

### v2

@@ 5.3 Path dependent option: the Call Sharpe option
Table 11: Call Sharpe. The value computed by [Guyon and Henry-Labordère, 2011] for the same option is $55.55$, while the benchmark obtained using a PDE approach is $58.40$. The column labeled “ME” gives the margin of error (95% CI half-width) for NNU results. Values in brackets indicate computational times (in seconds).
@@ 6 Conclusions
In this study, we demonstrated the efficacy of Machine Learning in enhancing the Uncertain Volatility Model for high-dimensional option pricing. Our findings indicate that both algorithms proposed – GTU and NNU – provide accurate and reliable results across a wide range of settings.
The NNU method proves particularly effective in high-dimensional problems, where its scalability and data-driven flexibility allow it to deliver accurate prices along with confidence intervals, which are valuable for risk assessment and regulatory reporting. Its ability to learn complex, non-linear relationships makes it especially suitable when the dimensionality the model increases.
In contrast, GTU remains more efficient in low-dimensional settings and in scenarios with uncertain correlation. Its tree-based structure ensures fast and reliable estimates with limited computational cost, making it an ideal choice when rapid evaluation is needed and the model dimensionality is moderate.
Overall, both methods contribute meaningfully to the development of data-driven superhedging strategies within the Uncertain Volatility Model framework. The integration of Machine Learning not only improves the precision of volatility and correlation control, but also enhances computational efficiency, providing robust tools for traders and risk managers. Future research may extend this work by investigating alternative learning architectures or hybrid approaches, further bridging the gap between theoretical models and practical implementations in quantitative finance.
Acknowledgement
This work was supported by the Departmental Strategic Plan (PSD) of the University of Udine. Interdepartmental Project on Artificial Intelligence (2020-2025).
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT (version 4o) in order to improve the grammatical quality and readability. After using this service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.